OpenStudy (anonymous):

The population of frogs in a park is decreasing exponentially at a rate of 6% each year. In the year 2007, the population was estimated to be 5500. a. Use the exponential decay formula to write an equation to model the situation. b. Use your equation from Part a to predict the frog population in the year 2018. c. Sketch the graph of the equation you wrote in Part a. Find its asymptote, domain, and range.

OpenStudy (anonymous):

y = a(1-b)^x y= final amount a=original amount x=time b= percent change(it would be + if it said increasing) note:change to decimal form. y=5500 x(1-6%)^(2018-2007) y= 5500(.94^11)